Developing and Using a Predetermined Overhead Rate: High-Low Cost Estimation
For years, Mattoon Components Company has used an actual plantwide overhead rate and based its prices on cost plus a markup of 30 percent. Recently the marketing manager, Holly Adams, and the production manager, Sue Walsh, confronted the controller with a common problem. The marketing manager expressed a concern that Mattoon's prices seem to vary widely throughout the year. According to Adams, "It seems irrational to charge higher prices when business is bad and lower prices when business is good. While we get a lot of business during high-volume months because we charge less than our competitors, it is a waste of time to even call on customers during low-volume months because we are raising prices while our competitors are lowering them." Walsh also believed that it was "folly to be so pushed that we have to pay overtime in some months and then lay employees off in others." She commented, "While there are natural variations in customer demand, the accounting system seems to amplify this variation."

(b.) Assume that the Mattoon Components Company had the following total manufacturing overhead costs and direct labor hours in 2016 and 2017:

2016 2017
Total manufacturing overhead $208,000 $250,000
Direct labor hours 20,000 28,000

Use the high-low method to develop a cost estimating equation for total manufacturing overhead.
Total costs = $Answer

+ $Answer

X

(c.) Develop a predetermined rate for 2018, assuming 25,000 direct labor hours are budgeted for 2018.
Round answers to two decimal places.
$Answer

per direct labor hour

(d.) Assume that the actual level of activity in 2018 was 30,000 direct labor hours and that the total 2018 manufacturing overhead was $250,000. Determine the underapplied or overapplied manufacturing overhead at the end of 2018.
Do not use a negative sign with your answer.
$Answer

AnswerUnderappliedOverapplied
Answer #1

b.

High low method :

Variable cost per unit = $250,000-208,000 / 28,000-20,000

Variable cost per unit = $42,000 / 8,000 = $5.25

Fixed cost per unit = Total cost - variable cost

Fixed cost per unit = $250,000 - (28,000*$5.25)

Fixed cost per unit = $250,000 - 147,000 = $103,000

Total cost = Fixed cost + variable cost per unit * X units

Total cost = $103,000 + $5.25 * X

c.

Predetermined overhead rate = Budgeted manufacturing overhead / Budgeted direct labor hours

Predetermined overhead rate = $103,000+($5.25*25,000) / 25,000

Predetermined overhead rate = $234,250 / 25,000 = $9.37 per direct labor hour

d.

Applied manufacturing overhead = (30,000*$5.25) + $103,000 = $260,500

Actual manufacturing overhead = $250,000

Overapplied overhead = $260,500 - 250,000 = $10,500